Sam scored 5 points more than Jasmine on a class test. If they scored a total of
153 points, what was Sam’s score?

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52781)

.
Sam scored 5 points more than Jasmine on a class test. If they scored a total of
153 points, what was Sam’s score?
~~~~~~~~~~~~~~~

Sam's score = x;

Jasmine's score = x-5.



Total scores equation is

    x + (x-5) = 153.


Simplify and find x

    2x - 5 = 153

    2x     = 153 + 5

    2x     =   158

     x     =   158/2 = 79.


ANSWER.  Sam's score is 79.

Solved and thoroughly explained.



Answer by Alan3354(69443)

Sam scored 5 points more than Jasmine on a class test. If they scored a total of
153 points, what was Sam’s score?
----------------
(153 + 5)/2 = 79